OpenStudy (anonymous):

Can someone check 6 parallel lines/angles - 8th grade.

jimthompson5910 (jim_thompson5910):

what do you need checked exactly?

OpenStudy (usman1995):

where is the question??

OpenStudy (anonymous):

Two interior angles of triangle PQR measure 27° and 79°. What is the measure of ∠PQR? A. 74° <--- B. 90° C. 174° D. 254° I'm not completely sure about this on....

OpenStudy (usman1995):

sum of all angles is 180 alright,, so 27+79+x=180 now subtract the sum of 27+79 from 180 and that's ur anwer

OpenStudy (usman1995):

u get it @CrazyCountryGirl

OpenStudy (anonymous):

So I'm right? :)

OpenStudy (usman1995):

answer is 74

jimthompson5910 (jim_thompson5910):

you are correct nice job

OpenStudy (usman1995):

thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

check: 27 + 79 + x = 180 27 + 79 + 74 = 180 180 = 180 ... everything checks out

OpenStudy (usman1995):

and my medal @CrazyCountryGirl plz